# Low-Latency Event-Based Velocimetry for Quadrotor Control in a Narrow Pipe

### v1

@@ VII-B2 Quantitative Results
Table III shows the results of our method and the selected baselines. We compare the RMSE (error here refers to the average endpoint distance of the flow vectors) on two sequences: the first is a slower sequence where the flow speeds are around 2 m/s and the quadrotor is sitting on the ground while spinning its propellers, and the second is a faster in-flight sequence with speeds up to 5 m/s. The real-time factor indicated in the table is $>1$ if the method runs faster than 500 Hz real-time. For contrast maximization, we do not report a real-time factor, as our implementation was not optimized for runtime.
TABLE III: Comparison of Optical Flow Methods on Smoke Data
@@
As Tab. III shows, the proposed template matching method quantitatively outperforms the contrast-maximization and the unsupervised method while achieving the highest real-time factor. This clearly shows that well-known optical-flow estimators are not necessarily ideal for smoke velocimetry applications.
@@ VII-C2 A Position-Only Model
TABLE IV: Model Comparison for Disturbance Estimation (RMSE)
@@ VII-C3 Position & Flow Model
Our full model presented in Sec. III-C utilizes both the ConvLSTM branch to process the flow measurements, as well as the position branch to account for the position of the quadrotor. This model is referred to as Pos. + Flow (RCNN). In addition to the recurrent ConvLSTM version, we also evaluate a non-recurrent variant where the ConvLSTM layers are replaced with standard 2D convolutions, labeled as Pos. + Flow (CNN). We evaluate the networks in terms of disturbance-wrench prediction RMSE on the test split, and the results are presented in Tab. IV.
Considering all test data, the estimation of the disturbance $f^{d}_{y}$ and $\tau^{d}_{x}$ improved by 38 % and 42 %, respectively, when also using the flow estimates for disturbance estimation, highlighting the usefulness of real-time smoke velocimetry in this scenario. For $f_{z}^{d}$, the improvement is only 4 %, indicating that potentially relevant flow structures might not be captured by our sparse velocimetry approach. However, considering that the $z$-direction on a quadrotor is directly actuated, and that a prediction error of 0.12 N RMS on a quadrotor requiring 2.14 N static hover thrust is only 6 %, this prediction accuracy is likely sufficient for control. We also observe that adding recurrency to the flow processing part of the network helps, as the RCNN is between 2 % and 6 % better in disturbance estimation compared to the CNN variant.
Next, we split the test dataset into steady-state sequences and dynamic sequences, depending on whether the quadrotor moved more than 5 cm laterally. From Tab. IV, we can see that the overall accuracy degrades slightly, but the difference between the position-only and the full input becomes more pronounced in dynamic sequences, with up to 59 % improvement in the roll disturbance. This highlights the importance of flow sensing in dynamic scenarios.
@@ VII-D1 Hovering Flight
To this end, we first analyze the performance of different disturbance-handling strategies during hovering flight, as summarized in Table V. We evaluate three policies trained under different configurations. The first, referred to as baseline, is trained without exposure to any aerodynamic disturbances and with zeroed-out disturbance observations. The two other policies, wo/ dist. obs. and w/ dist. obs., are trained with the full aerodynamic disturbance simulation as described in Sec. V-B. The difference between the two is that the wo/ dist. obs. policy has zeroed-out disturbance observations during training. The policy w/ dist. obs. (ours) also has access to real-time information about the disturbance.
In Tab. V, the standard deviation and inter-quartile ranges of the hover position are shown. We observe that incorporating disturbance modeling during training improves hover accuracy, as both the wo/ dist. obs. and w/ dist. obs. policies perform much better than the baseline. This improvement is likely due to the consistent aerodynamic wrenches present in the confined environment, which tend to push the quadrotor outward from the centerline. A controller trained under such conditions learns to anticipate and compensate for this instability. Furthermore, comparing the second and third policies, we find that access to real-time disturbance observations further enhances control performance. Specifically, our policy improves position standard deviation in the $y$-direction by 29 %, demonstrating the value of integrating real-time event-based flow information into the control pipeline.
TABLE V: Hover Position Deviation (Flight in Center)
@@ VII-D2 Lateral Translation
While improvements during hover are observable, the advantages of flow-aware disturbance estimation become even more apparent in dynamic flight scenarios. To investigate this in detail, we analyze a lateral translation maneuver in which the quadrotor is commanded to move from 50 mm to -40 mm in the $y$-direction. The resulting position trajectories in $y$ and $z$ of one experiment are shown exemplarily in Fig. 9. During this maneuver, the quadrotor must remain within ±$\pm 110$ mm in the $y$-direction to avoid crashing into the pipe wall.
@@ VIII-B Limitations
Lastly, we want to discuss the choice of an event camera instead of a more standard high-speed camera for flow sensing. At the same resolution and a flow speeds of 5 m/s, we would need a 1/2500–1/4000s exposure time to limit motion blur to below 2 px. Achieving sufficient illumination for high-speed standard imaging inside the pipe would ideally require a pulsed laser sheet, which is impractical due to safety and engineering constraints. When using our bright LED setup instead, high-speed video would require very high-end sensors to get acceptable noise levels and result in an expensive, bulky setup. Furthermore, most high-speed cameras are not optimized for machine vision and do not provide a low-latency image stream as output. Event cameras, on the other hand, provide high temporal resolution and low latency under the same conditions, at significantly lower cost and complexity.
@@ IX Conclusion
To the best of our knowledge, this is the first time that closed-loop control with real-time flow field feedback has been demonstrated on an aerial robot. We believe this opens new directions for research on robust flight in turbulent and confined environments where standard sensing and control approaches are insufficient.
@@ A-A Choice of LEDs
In this work, the Osram SFH4350 [71] IR-LED 3 mm was selected, as it features an extremely short switching time of 12 ns. Being designed for optoelectronics, it also permits up to 1 A pulsed forward current for pulses shorter than 100 $\mu$s, if the duty cycle is below 2 %. The LED has its peak emission around 850 nm, with a spread of 50 nm, which is within the sensitivity range of most CMOS-based imaging sensors.
@@ A-B Circuit Design
    If the LEDs blink too fast, measuring the signal becomes difficult. Typically, event cameras perform very well at measuring signals with frequencies up to 2-3 kHz [72, 73, 74], and the events are timestamped with 1 $\mu$s time resolution.
@@
TABLE VI: Resistor and capacitor values for the different LEDs. The calculated periods as well as the measured frequencies $f$ and duty cycles $\alpha$ are listed.
@@
Figure 10: Current through an LED blinking at the lowest frequency of 1.73 kHz. The upper plot shows the entire period $t=t_{\text{on}}+t_{\text{off}}$, whereas the lower plot shows only the pulse during which the LED is on for 3.8 $\mu$s. From the plot, we get a switch-on time constant $\tau_{\text{on}}$ of the LED of 84 ns (time to reach 63 % of the steady-state value).
@@
The resistor and capacitor values used in the NE555 timer circuit are listed in Tab. VI. The values have been calculated such that the frequencies of all 5 LEDs satisfy points 2 to 5. After building and manufacturing the PCB, the measured frequencies are also listed in Tab. VI. Note that the mismatch with respect to the calculated values is about 20 %. This mismatch is consistent across 5 identical copies of the board and is of no concern for practical applications, as it is straightforward to measure the blinking frequency with an oscilloscope. Exemplarily, Fig. 10 shows the current through one LED.

### v2

@@ VII-B2 Quantitative Results
Table V shows the results of our method and the selected baselines. We compare the RMSE (error here refers to the average endpoint distance of the flow vectors) on two sequences: the first is a slower sequence where the flow speeds are around 2 m/s and the quadrotor is sitting on the ground while spinning its propellers, and the second is a faster in-flight sequence with speeds up to 5 m/s. The real-time factor indicated in the table is >1>1 if the method runs faster than 500 Hz real-time. For contrast maximization, we do not report a real-time factor, as our implementation was not optimized for runtime.
TABLE V: Comparison of Optical Flow Methods on Smoke Data
@@
As Tab. V shows, the proposed template matching method quantitatively outperforms the contrast-maximization and the unsupervised method while achieving the highest real-time factor. This clearly shows that well-known optical-flow estimators are not necessarily ideal for smoke velocimetry applications.
@@ VII-C2 A Position-Only Model
TABLE VI: Model Comparison for Disturbance Estimation (RMSE)
@@ VII-C3 Position & Flow Model
Our full model presented in Sec. III-C utilizes both the ConvLSTM branch to process the flow measurements, as well as the position branch to account for the position of the quadrotor. This model is referred to as Pos. + Flow (RCNN). In addition to the recurrent ConvLSTM version, we also evaluate a non-recurrent variant where the ConvLSTM layers are replaced with standard 2D convolutions, labeled as Pos. + Flow (CNN). We evaluate the networks in terms of disturbance-wrench prediction RMSE on the test split, and the results are presented in Tab. VI.
Considering all test data, the estimation of the disturbance $f^{d}_{y}$ and $\tau^{d}_{x}$ degrades by 38 % and 42 %, respectively, when not using the flow estimates for disturbance estimation, highlighting the usefulness of real-time smoke velocimetry in this scenario. For $f_{z}^{d}$, the improvement is only 4 %, indicating that potentially relevant flow structures might not be captured by our sparse velocimetry approach. However, considering that the $z$-direction on a quadrotor is directly actuated, and that a prediction error of 0.12 N RMS on a quadrotor requiring 2.14 N static hover thrust is only 6 %, this prediction accuracy is likely sufficient for control. We also observe that adding recurrency to the flow processing part of the network helps, as the RCNN is between 2 % and 6 % better in disturbance estimation compared to the CNN variant.
Next, we split the test dataset into steady-state sequences and dynamic sequences, depending on whether the quadrotor moved more than 5 cm laterally. From Tab. VI, we can see that the overall accuracy degrades slightly, but the difference between the position-only and the full input becomes more pronounced in dynamic sequences, with up to 59 % difference in the roll disturbance. This highlights the importance of flow sensing in dynamic scenarios.
Figure 10: Comparison of the average normalized errors of the disturbance estimator under increasing dropouts of flow measurements. The error is normalized w.r.t. the error of the Position model and averaged across $f_{y}$, $f_{z}$, and $\tau_{x}$. Our RCNN architecture maintains its performance up to 25 % dropouts of flow measurements.
Finally, we analyze the performance of the disturbance estimator under flow measurement dropouts with dropout probability $\alpha$ focusing on the case of completely dropped (e.g. zeroed) measurements across entire frames. Figure 10 shows the performance of the networks using flow information relative to the position-only network. We see that, in the absence of flow measurements (i.e. $\alpha=1$), the performance of the networks is marginally worse compared to position-only. Our recursive RCNN architecture shows good robustness against dropouts, with the accuracy being practically unaffected until 25 % random dropouts, whereas the non-recurrent CNN immediately shows a decrease in performance with dropped measurements.
@@ VII-D1 Hovering Flight
To this end, we first analyze the performance of different disturbance-handling strategies during hovering flight, as summarized in Table VII. We evaluate three policies trained under different configurations. The first, referred to as baseline, is trained without exposure to any aerodynamic disturbances and with zeroed-out disturbance observations. The two other policies, wo/ dist. obs. and w/ dist. obs., are trained with the full aerodynamic disturbance simulation as described in Sec. V-B. The difference between the two is that the wo/ dist. obs. policy has zeroed-out disturbance observations during training. The policy w/ dist. obs. (ours) also has access to real-time information about the disturbance during training.
In Tab. VII, the standard deviation and inter-quartile ranges of the hover position are shown. We observe that incorporating disturbance modeling during training improves hover accuracy, as both the wo/ dist. obs. and w/ dist. obs. policies perform much better than the baseline. This improvement is likely due to the consistent aerodynamic wrenches present in the confined environment, which tend to push the quadrotor outward from the centerline. A controller trained under such conditions learns to anticipate and compensate for this instability. Furthermore, comparing the second and third policies, we find that access to real-time disturbance observations further enhances control performance. Specifically, our policy improves position standard deviation in the $y$-direction by 29 %, demonstrating the value of integrating real-time event-based flow information into the control pipeline. As described in Sec. V-A we simulate measurement dropouts during training. Our experiments empirically validate this approach: if no smoke is present in the pipe the w/ dist. obs. (ours) policy has no real-time disturbance observations and it performs on par with its wo/ dist. obs. counterpart.
TABLE VII: Hover Position Deviation (Flight in Center)
@@ VII-D2 Lateral Translation
While improvements during hover are observable, the advantages of flow-aware disturbance estimation become even more apparent in dynamic flight scenarios. To investigate this in detail, we analyze a lateral translation maneuver in which the quadrotor is commanded to move from 50 mm to -40 mm in the $y$-direction. The resulting position trajectories in $y$ and $z$ of one experiment are shown exemplarily in Fig. 11. During this maneuver, the quadrotor must remain within $\pm 110$ mm in the $y$-direction to avoid crashing into the pipe wall.
@@ VIII-B Limitations
Lastly, we want to discuss the choice of an event camera instead of a more standard high-speed camera for flow sensing. At the same resolution and a flow speeds of 5 m/s, we would need a 1/2500–1/4000s exposure time to limit motion blur to below 2 px. To assess the feasibility, we first calculate the luminance of the scene based on the photos [73]:
$L=(A^{2}/T)\cdot(K/S)=3.5\,\mathrm{cd/m^{2}}$ (12)
where $L$ is the luminance, $A$ the aperture (f-stop number), $T$ the exposure time, $S$ the ISO sensitivity, and $K$ a calibration factor in the range of 10.6 to 13.4 [73]. We use $K=12.5\,\mathrm{cd\,s/m^{2}}$ and obtain a fairly low luminance on the order of $10^{0}$. This luminance level already requires twilight-vision in humans [74]. To achieve the shutter speeds around 1/2500-1/4000 in a conventional high-speed camera, this would amount to ISO 25600 with an f/1.4 lens. Only the most advanced high-speed cameras can even achieve such high sensitivity, but the image quality will be severely degraded due to noise.
Therefore, achieving sufficient illumination for high-speed standard imaging inside the pipe would ideally require a pulsed laser sheet, which is impractical due to safety and engineering constraints. Furthermore, most high-speed cameras are not optimized for machine vision and do not provide a low-latency image stream as output. Event cameras, on the other hand, provide high temporal resolution and low latency under the same conditions, at significantly lower cost and complexity.
The sparse event-stream also provides significant advantages in terms of computational load compared to frame-based high-speed video cameras. This enables running our event-based motion-capture on small embedded computers such as the Raspberry Pi 4B. Similarly, the event-based smoke velocimetry can be deployed on Nvidia’s Jetson family, enabling future applications where onboard deployments are required. Designing a larger drone platform which tightly integrates the light sheet and smoke generation system, a lightweight event camera and a Jetson computer into a single platform would be a next step allowing real-time flow sensing to be performed in real-world environments without relying on an instrumented pipe.
@@ IX Conclusion
To the best of our knowledge, this is the first time that closed-loop control with real-time flow field feedback has been demonstrated on an aerial robot. We believe this opens new directions for research on robust flight in turbulent and confined environments where standard sensing and control approaches are insufficient. Additionally, understanding how characteristic flow patterns and disturbance magnitudes depend on the pipe’s cross-section and the drone’s geometry and mass offers an exciting avenue for future fluid dynamics research.
@@ A-A Choice of LEDs
[…]
@@ A-B Circuit Design
[…]
@@
TABLE VIII: Resistor and capacitor values for the different LEDs. The calculated periods as well as the measured frequencies $f$ and duty cycles $\alpha$ are listed.
@@
[…]
@@
The resistor and capacitor values used in the NE555 timer circuit are listed in Tab. VIII. The values have been calculated such that the frequencies of all 5 LEDs satisfy points 2 to 5. After building and manufacturing the PCB, the measured frequencies are also listed in Tab. VIII. Note that the mismatch with respect to the calculated values is about 20 %. This mismatch is consistent across 5 identical copies of the board and is of no concern for practical applications, as it is straightforward to measure the blinking frequency with an oscilloscope. Exemplarily, Fig. 12 shows the current through one LED.
@@ Appendix B Pose Estimation Noise
Figure 14: Effects of marker occlusion on the SqPnP accuracy. The drone is placed at different positions along the pipe (pipe center is at a distance of 3 m). Comparing the pose estimation error between the occluded and the unoccluded case we see that the effects of occlusion are very well handled by SqPnP.
In addition to the above comparison between EPnP and SqPnP, we also analyze the performance of SqPnP when a single marker is occluded and thus only $N=4$ markers are visible. The experiments are conducted inside the pipe with the Zeiss 50 mm lens under identical conditions to those used for flight experiments. Fig. 14 shows the longitudinal ($z_{\mathcal{C}}$) and lateral position error as well as the average attitude error. Overall, the difference between the 4 marker and 5 marker case is small and, from a practical perspective, not relevant. Interestingly, in the longitudinal direction we observe that using 4 vs. 5 markers strictly increases the measurement noise, whereas at close distances the measurement noise parallel to the image plane and in the attitude is slightly lower for 4 markers. This is counterintuitive at first, but could be explained if we consider imperfect 3D positions of the LEDs on the drone. With 5 markers the solver has more ability to ’jump’ between solutions, increasing the noise level.